A woman is 10 years older than her brother. Next year she will be twice as old as him? How old are they now? How do I write this as an equation?
Answer by mananth(16946)

let brother's age be x
her age will be x+10
..
ages after 1 year
brother =x+1
woman = x+11
..
x+11=2(x+1)
x+11-2x+2
2x-x=11-2
x=9 brother's age
her age = x+10= 9+10=19 years